OpenStudy (hihi67):

please solve:5k-(m+4k-(3m-2k)+5m)-k

OpenStudy (hihi67):

5k-(m+4k-(3m-2k)+5m)-k 5k-(m+4k-(3m-2k)+5m)-k

OpenStudy (anonymous):

hello!

OpenStudy (hihi67):

Hello!

OpenStudy (anonymous):

5k-(m+4k-(3m-2k)+5m)-k =5k -(m+4k -3m+2k +5m) -k =5k -m-4k+3m-2k-5m-k =-2k -3m Take note when expanding the brackets. Step by step.

OpenStudy (hihi67):

i get the whole expanding brackets thing. i have the entire thing with the brackets right! the in between part, thats where i got it wrong